a tortoise makes a journey in two parts; it can either walk at 4cm/s or crawl at 3cm/s. if the tortoise walks the first part and crawls the second, it takes 110 seconds. If it crawls the first part and walks the second, it takes 100 seconds. find the lengths of the two parts of the journey.
Answer by mananth(16949)

let it walk x cm
let it crawl y cm
..
x/4 + y/3 = 110
3x+4y /12= 110
3x+4y = 1320.....................1
...
x/3 + y/4 = 100
4x+3y / 12 = 100
4x+3y= 1200......................2
...
multiply equation 1 by -4 & equation 2 by 3
-12x-16y= -5280
12x+9y = 3600
...
add the two equations
...
-7 y = -5280+3600
-7y = -1680
y = -1680/-7
y = 240 cm crawl
plug value of y in equation 1
3x+4y=12*110
3x+960=1320
3x=1320-960
3x=360
x = 120 cm walking

m.ananth@hotmail.ca